# Application of Long-Short Term Memory and Convolutional Neural Networks for Real-Time Bridge Scour Forecast

### v1

@@ 5.2.1 Scour Forecast for Alaska Bridges
Fig.  8 and Fig.  9 present the predicted values of Sonar (bed elevation) and Stage (water elevation) by the top-5 models for bridge $230$ and $539$ across the test dataset, between July-$2021$ to Sep-$2021$. The performance of models is assessed based on bed elevation (scour depth) predictions.
@@ 5.2.2 Scour Forecast for Oregon Bridges
Stage and Sonar prediction of top-5 models over the test dataset is presented in Fig.  10 and Fig.  11 for Luckiamute and Trask bridges, respectively. The test data is selected for the period between April-$2021$ to July-$2021$ and August-$2021$ to Nov-$2021$, for Luckiamute and Trask, where significant scour and fillings are observed.
@@ 5.2.3 Sequential Training on Oregon Data
Table 4 presents the MAE values on Sonar, and Fig.  13 shows the time-series forecast over the final-test data using the 3-Fold and 5-Fold sequentially-trained LSTM models for Trask bridge. It is observed that the prediction error is reduced using the Sequential Training method.
@@ 5.3 CNN Evaluation (Experiment 3)
5.3.1 Scour Forecast for Alaska Bridges
Fig.  14 and Fig.  15 demonstrate the time-series prediction of Sonar and Stage by the top-5 FCN models across the test dataset. The CNN models demonstrate a reasonable performance, comparable with LSTM models, capturing the trend of scour and filing a week in advance.
@@ 5.3.1 Scour Forecast for Alaska Bridges
The CNN model with full convolution from Equation  14 shows that it outperforms other convolutional strategies in generalization and captures both Sonar and Stage trends. Particularly, DCN models fail to capture the scour and filling trends as compared to the VCN and FCN. This implies that skipping the intermediate observations by the dilated convolutions from Equation  16 did not result in higher accuracy. One reason behind the superior performance by FCN can be the inclusion of the edging values while making batch-to-batch transitions via the 00-$padded$ convolution. This ensures no loss of information while learning the underlying scour patterns.
@@ 5.3.2 Scour Forecast for Oregon Results
Fig.  16 illustrates the performance of top-5 FCN models over the test dataset, between April 2021 and July 2021. Similar to the LSTM models, the CNN models perform less accurately for Oregon data compared with Alaska, particularly in capturing the scour and fill trends. The error of mean Sonar prediction varies between 1.5 to 2.5ft (0.5 to 0.75m).
@@ 5.4 Feature Combination Analysis (Experiment 4)
Different combinations of features are considered as input to LSTM models to evaluate their impact in scour prediction (wrapper method). The best Single-Shot LSTM configurations were trained for this purpose over two Alaska and Oregon bridges as shown in Figs. 17 and 18. Note that for Oregon bridges, the discharge data was not; therefore, dC dV are not included in the impact analysis.
@@ 6 Conclusions
This paper investigated various prominent DL algorithms for scour forecast based on historic sensor monitoring data, including three variants of LSTM and CNN, for case-study bridges in Alaska and Oregon. The primary focus was to address the challenges of DL models for real-time scour prediction, including performance optimization of algorithms and hyperparameter tuning in terms of both accuracy and computational efficiency, generalisation to diverse geological, geomorphological and hydraulic/flow characteristics, and impact of various combinations of sensor features relevant to scour.
The top-performing LSTM models achieved mean absolute error (MAE) ranging from 0.1m (in Alaska) to 0.5m (in Oregon). The best CNN model, the FCN variant, showed similar error ranges (0.1m to 0.75m). Both Deep Learning (DL) algorithms showed promising performance in forecasting scour for bridges in Alaska, with CNN models, particularly the FCN variant, achieving slightly lower prediction errors. However, CNN was trained with significantly lower computational costs running on the same GPU clusters. Higher errors were observed for Oregon bridges compared to Alaska, especially in capturing peaks of scour and filling episodes. This was related to the fundamentally different scour process in Oregon bridges, such as tidal flows and the presence of coastal bed forms, and the lack of substantial historical data. In addition, Oregon bridges did not show consistent periodical and seasonal flow and scour patterns as observed in Alaskan bridges.

### v2

@@ 5.2.1 Scour Forecast for Alaska Bridges
Fig.  9 and Fig.  10 present the predicted values of Sonar (bed elevation) and Stage (water elevation) by the top-5 models for bridge $230$ and $539$ across the test dataset, between July-$2021$ to Sep-$2021$. The performance of models is assessed based on bed elevation (scour depth) predictions.
@@ 5.2.2 Scour Forecast for Oregon Bridges
Stage and Sonar prediction of top-5 models over the test dataset is presented in Fig.  11 and Fig.  12 for Luckiamute and Trask bridges, respectively. The test data is selected for the period between April-$2021$ to July-$2021$ and August-$2021$ to Nov-$2021$, for Luckiamute and Trask, where significant scour and fillings are observed.
@@ 5.2.3 Sequential Training on Oregon Data
Table 3 presents the MAE values on Sonar, and Fig. 14 shows the time-series forecast over the final-test data using the 3-Fold and 5-Fold sequentially-trained LSTM models for Trask bridge. It is observed that the prediction error is reduced using the Sequential Training method.
@@ 5.3 CNN Evaluation (Experiment 3)
5.3.1 Scour Forecast for Alaska Bridges
Fig.  15 and Fig.  16 demonstrate the time-series prediction of Sonar and Stage by the top-5 FCN models across the test dataset. The CNN models demonstrate a reasonable performance, comparable with LSTM models, capturing the trend of scour and filing a week in advance.
The CNN model with full convolution from Equation 14 shows that it outperforms other convolutional strategies in generalization and captures both Sonar and Stage trends. Particularly, DCN models fail to capture the scour and filling trends as compared to the VCN and FCN. This implies that skipping the intermediate observations by the dilated convolutions from Equation  16 did not result in higher accuracy. One reason behind the superior performance by FCN can be the inclusion of the edging values while making batch-to-batch transitions via the 00-$padded$ convolution. This ensures no loss of information while learning the underlying scour patterns.
@@ 5.3.2 Scour Forecast for Oregon Results
Fig.  17 illustrates the performance of top-5 FCN models over the test dataset, between April 2021 and July 2021. Similar to the LSTM models, the CNN models perform less accurately for Oregon data compared with Alaska, particularly in capturing the scour and fill trends. The error of mean Sonar prediction varies between 1.5 to 2.5ft (0.5 to 0.75m).
@@ 5.4 Feature Combination Analysis (Experiment 4)
Different combinations of features are considered as input to LSTM models to evaluate their impact in scour prediction (wrapper method). The best Single-Shot LSTM configurations were trained for this purpose over two Alaska and Oregon bridges as shown in Figs. 18 and 19. Note that for Oregon bridges, the discharge data was not; therefore, dC dV are not included in the impact analysis.
@@ 6 Conclusions
This paper investigated prominent Deep Learning (DL) algorithms for scour forecasting based on historic sensor monitoring data, including three variants of LSTM and CNN, for case-study bridges in Alaska and Oregon. The primary focus was to address the challenges of DL models for real-time scour prediction, including 1) performance optimization of algorithms and hyperparameter tuning, 2) generalisation to diverse locations with distinct geological/geomorphological and hydraulic/flow characteristics, and 3) impact of various combinations of sensor features relevant to scour.
The top-performing LSTM models achieved mean absolute error (MAE) ranging from 0.1m (in Alaska) to 0.5m (in Oregon). The best CNN model, the FCN variant, showed similar error ranges (0.1m to 0.75m). Both DL algorithms showed promising performance in forecasting scour for bridges in Alaska, with CNN models, particularly the FCN variant, achieving slightly lower prediction errors. However, CNN was trained with significantly lower computational costs running on the same GPU clusters. Higher errors were observed for Oregon bridges compared to Alaska, especially in capturing peaks of scour and filling episodes. This was related to the fundamentally different scour process in Oregon bridges, such as tidal flows and the presence of coastal bed forms, and the lack of substantial historical data. In addition, Oregon bridges did not show consistent periodical and seasonal flow and scour patterns as observed in Alaskan bridges.